OpenStudy (anonymous):

Algebra II Question! Solve for k, given points (1, k) and (5, | k |) lie on a line that has a slope of 3/5 . A) -3/10 B) -6/5 C) -3/20 D) Ø E) NOTA

OpenStudy (anonymous):

I would use vector geometry, with the given slope.

OpenStudy (anonymous):

It's Algebra II :P

OpenStudy (anonymous):

it's not in the title q-:

OpenStudy (anonymous):

Now it is :P XD

OpenStudy (anonymous):

*grins*

OpenStudy (kinggeorge):

I would use the fact that since it has a slope that isn't infinity, \(|k|\neq k\). Hence, \(k\) must be negative.

OpenStudy (anonymous):

Oh that makes sense...but all of the answer choices are negative, so I kind of assumed that...

OpenStudy (kinggeorge):

Next up, that means that \[\frac{-k-k}{5-1}=\frac{-2k}{4}=\frac{3}{5}\]Solve for \(k\), and I get \[k=-\frac{6}{5}\]

OpenStudy (kinggeorge):

I used \((-k)-k\) for the numerator since \(|k|=-k\) must be true since \(k\) is negative.

OpenStudy (anonymous):

Oh. I get it now...I just for some reason tend to get confused when absolute value comes in...

OpenStudy (anonymous):

Thank you :)

OpenStudy (kinggeorge):

You're welcome.